# Yesnt: Are Diffusion Relighting Models Ready for Capture Stage Compositing? A Hybrid Alternative to Bridge the Gap

### v1

@@ Abstract
Figure 1: We integrate diffusion-based decomposition priors with variational methods and image-based lighting to achieve physically plausible, temporally stable relighting of captured volumetric content and seamless compositing into diverse virtual environments
@@ 1 Introduction
The first challenge, novel view synthesis (NVS), has witnessed extraordinary progress in recent years. Breakthroughs such as Neural Radiance Fields (NeRF) [52] and 3D Gaussian Splatting [36], together with a large body of follow-up work [17], have demonstrated the ability to generate photorealistic renderings from sparse views with remarkable fidelity. By combining multi-view cues with expressive scene representations that capture geometry and appearance, these approaches have transitioned from academic prototypes to production-ready tools, establishing a new foundation for volumetric video pipelines [69, 1].
@@
The second challenge, relighting, has been extensively studied, but remains difficult in capture stages where inputs are sparse, lighting is uniform, and material or illumination cues are limited. While one-light-at-a-time (OLAT) capture can enable high-quality relighting [73, 66], it remains impractical for most applications. Inverse rendering provides a principled formulation, yet becomes ill-posed under limited observations needed for dynamic acquisition. Recent data-driven approaches, including diffusion-based decomposition models, offer a promising alternative by directly predicting material and illumination properties from images [58, 81, 34, 43].
This motivates our central question: Are diffusion relighting models ready for capture stage compositing? While their progress is promising, it remains unclear whether these models can meet the demands of real-world use when integrated into volumetric capture pipelines. In this work, we revisit the relighting problem from a practical perspective with an emphasis on stability. We propose a simple yet effective approach that leverages priors from state-of-the-art decomposition models while avoiding the complexity and instability of full inverse rendering or generative diffusion methods. To this end, we introduce a flow-guided temporal regularization scheme for consistent material estimates and a lightweight screen-space relighting module inspired by real-time graphics pipelines with efficient shadow computation. By combining learned priors with physically informed design choices, our method achieves more stable and accurate relighting, making it better suited for production pipelines in volumetric capture.
Contributions.
In summary, our contributions are
@@
  • An end-to-end relighting pipeline for volumetric capture, where we present a practical, production-ready relighting framework that integrates state-of-the-art image decomposition priors with physically-motivated rendering, enabling consistent and photorealistic results across dynamic sequences.
  • Temporal regularization for dynamic relighting that mitigates flickering and appearance instability, preserving both fine details and smooth lighting transitions over time.
@@
  • Comprehensive evaluation of the usability of diffusion-based relighting approaches for in real-world volumetric capture scenarios, highlighting trade-offs in quality, stability, and computational cost.
@@ 2 Related Work
Material estimation and relighting aim to recover surface reflectance properties and scene illumination, enabling realistic appearance synthesis under novel lighting conditions. Classical approaches formulate this task as an inverse rendering problem; owing to its inherent ill-posedness, early methods rely on strong, hand-crafted priors [5, 4, 10, 84]. A closely related direction focuses on estimating material appearance parameterized as a spatially varying BRDF (SVBRDF), again leveraging carefully designed priors [49, 2]. The advent of deep learning has relaxed many of these constraints, enabling SVBRDF estimation from a single texture image [18, 67] or from multi-view observations [19]. More recently, deep learning-based inverse rendering frameworks have coupled differentiable rendering with neural networks to jointly recover geometry, reflectance, and illumination from single-view [76, 39, 40, 61, 72], dual-view [7], or multi-view inputs [16, 46].
Recent years have seen significant progress in neural methods for relighting, which aim to synthesize novel illuminations for existing scenes or objects. Neural Radiance Fields (NeRF) [52] have been extended to support relighting by explicitly modeling material and illumination properties, via intrinsic decomposition of scene properties [8, 83, 60, 64], or via a learned scene representation using One-Light-at-a-Time (OLAT) recordings [73]. The introduction of 3D Gaussian Splatting (3DGS) [36] has led to diverse relighting strategies, via intrinsic decomposition [28, 22, 45, 62, 44], or relighting using OLAT data [20, 6, 48]. Some neural relighting methods avoid 3D reconstruction altogether, operating entirely in image space, e.g. by modelling view- and light-dependent transport as a learned function [82] or training a CNN [74] to predict relit results from sparse lighting samples.
More recently, diffusion models with their strong generative expression have been applied to relighting problems. Among these, Ren et al. proposed an image hormization, where they focus on light adaptation based on the background [58], followed by IC-Light [81] addressing illumination harmonization via text prompt guided background generation. Various works discuss relighting via text prompts [38, 34, 80] or light controlled image generation [79]. Neural Gaffer [34] propose relighting by either text prompt or environment map. While these approaches provide good results for single view images, the state-of-the-art Diffusion Renderer [43] unifies inverse and forward rendering within a video diffusion framework. This approach can reduce temporal inconsistencies – which are inherent in single-image settings, since diffusion models tend to produce different outputs for even slightly different inputs – but its applicability is limited by the large memory requirements. These constraints make it less suitable for high-resolution, multi-view capture stages, where consistency across views is essential. We discuss potential strategies to make its use feasible in such scenarios.
@@
Novel-view synthesis aims to generate images of a scene from viewpoints not present in the input observations. Early approaches for novel view synthesis primarily relied on rendering from images, and depth-guided warping techniques [23, 31, 12, 13]. Learning-based methods have since been proposed for novel view synthesis using multiplane images [25, 41, 51, 63] or voxel embeddings [33, 50]. More recently, neural radiance fields became the prevalent approach for novel view synthesis [52]. Since then, many variants focusing on more efficient scene representation have been proposed, such as low-rank factorized 4D grid scene representation [14], sparse 3D voxel grids with spherical harmonics [26], additional non-linearity in grid values [35], and directly optimized dense voxel grids [65]. Beyond static scene acceleration, research has expanded NeRFs to dynamic environments [57, 55, 56, 42, 27], enabling relighting by scene decomposition [8, 83, 9, 64]; improving the rendering of glossy surfaces via alternative radiance parameterizations [68], developing face-specific relighting methods [60], while others proposed scene editing without re-training [78]. Real-time rendering has likewise become a key focus [53, 75, 29], and recent work has moved toward alternative representations such as 3D Gaussian Splatting [36], which replaces implicit fields with explicit point-based primitives for real-time rendering. Subsequent work has further improved both rendering speed [32, 3, 24, 30, 54] and quality [70, 47, 71]. Additional advances have extended Gaussian-based methods beyond basic rendering, such as incorporating subsurface scattering [21] and enabling real-time relighting of head avatars [59]. In parallel, researchers have explored Gaussian-based surface reconstruction techniques [77, 15], further broadening their applicability to geometry recovery tasks.
@@ 3 Hybrid Relighting Framework
Our method combines three key components as shown in Fig. 2: volumetric scene representations for novel view synthesis and proxy geometry (Sec. 3.1), estimation and regularization of intrinsic material properties (Sec. 3.2 and Sec. 3.3) and a physically based relighting stage that produces view-consistent shading under novel illumination (Sec. 3.4). The final compositing step accounts for indirect interactions such as cast shadows and reflections, ensuring that the subject integrates plausibly with its surroundings. In the following, we describe each of these stages in detail.
@@ 3.1 Novel View Synthesis & Indirect Interactions
Given a multi-camera capture of a performance, our goal is to render the subject from novel viewpoints and to integrate it into a new environment with plausible indirect interactions, specifically cast shadows on nearby receivers and reflection contributions, while supporting relighting. To achieve this, we ideally need a NVS representation that provides (i) RGB appearance, (ii) depth maps, (iii) surface normals, and (iv) a proxy geometry suitable for integration into conventional rendering engines. Gaussian Opacity Fields (GOF) [77] are a natural choice, as they jointly optimize for appearance and normals, produce accurate depth estimates, and allow mesh extraction for downstream use.
However, we observe that GOF reconstructions often suffer from artifacts: curved, natural surfaces are approximated as piecewise planar surfaces, and spurious holes may appear. To mitigate these issues, we introduce two additional regularization terms during optimization: a Huber-regularized total variation loss on the rendered depth map, and an edge-aware Laplacian smoothness prior on the Gaussian normal map. Let $z(x,y)$ denote the rendered depth map. We enforce spatial smoothness with a Huber loss–based total variation penalty:
$\mathcal{L}_{\text{depth}}=\sum_{x,y}H_{\delta}(\tfrac{\partial z}{\partial x})+H_{\delta}(\tfrac{\partial z}{\partial y})$ (1)
where $H_{\delta}$ denotes Huber loss (cf. Eq. 6). To prevent blocky surface approximations while respecting appearance discontinuities, we penalize the Laplacian of the estimated normal field $n(x,y)$ in an edge-aware fashion with respect to the RGB image $I(x,y)$ as
@@
The extracted GOF mesh enables conventional rendering of indirect effects in external engines such as Blender. By inserting the mesh into the background scene and disabling intersections with foreground camera rays, we obtain plausible cast shadows and reflection contributions without requiring perfect geometry or material accuracy. Since indirect effects are typically low-frequency, coarse geometry suffices for the majority of capture scenarios to render plausible background images.
3.2 Hybrid G-Buffer Estimation
We estimate G-buffer channels using learned predictions for view- and light-independent material properties. We consider the problem of predicting roughness $r$ and metallic $\mu$ maps, from a single color image $I\in\mathbb{R}^{H\times W\times 3}$ by employing the intrinsic decomposition model $\mathcal{D}_{\theta}$ from [43] that provides the mapping
$\mathcal{D}_{\theta}:I\rightarrow\{r,\mu\}.$ (3)
@@ 3.2 Hybrid G-Buffer Estimation
where $\xi_{i}$ denotes the diffusion noise. These patchwise predictions are then blended into full-resolution maps $\tilde{x}^{(i)}$, yielding an ensemble $\{\tilde{x}^{(1)},\dots,\tilde{x}^{(K)}\}$ from which we compute per-pixel means and variances, thereby suppressing sampling variability and exposing an uncertainty signal useful for downstream processing.
@@ 3.3 Temporal Consistency via Optical Flow-Guided Regularization
We next enforce consistency across neighboring renderings. Let $\tilde{x}\in\{\tilde{r},\tilde{\mu}\}$ denote the per-pixel mean prediction and $\sigma_{x}$ the corresponding normalized standard deviation. Let $m_{\alpha}=\mathbbm{1}_{\alpha>0.5}$ be a binary foreground mask, and let $\mathcal{W}$ denote backward warping from frame $t+1$ to $t$ using estimated optical flow. We write $\odot$ for element-wise multiplication and $\mathcal{V}$ to denote a binary map of valid flows which point to valid foreground $\mathcal{W}(m_{\alpha})$ and are forward/backward consistent. The temporally regularized estimate $x_{t}$ at step $t$ is obtained by minimizing
$\begin{split}\underset{x}{\arg\min}&\ \sum_{t}\langle H_{\delta}(\tilde{x}_{t}-x_{t}),1-\sigma_{x,t}\rangle+\lambda_{1}TV_{\varepsilon}(x_{t})\\ &+\lambda_{2}\left\|\left(\mathcal{W}(x_{t})-x_{t}\right)\odot\mathcal{V}\right\|_{2}^{2}\end{split}$ (5)
where $\langle\cdot,\cdot\rangle$ is the element-wise inner product and $H_{\delta}$ is the (element-wise) Huber loss
$H_{\delta}(a)=\begin{cases}\frac{1}{2}a^{2}&\text{if }|a|<\delta\\ \delta\left(|a|-\frac{1}{2}\delta\right)&\text{otherwise.}\end{cases}$ (6)
The (smoothed) isotropic total variation is
$\mathrm{TV}_{\varepsilon}(x^{t})\;=\;\sum_{i}\sqrt{\|\nabla_{x}x^{t}_{i}\|_{2}^{2}+\|\nabla_{y}x^{t}_{i}\|_{2}^{2}+\varepsilon},$ (7)
for small $\varepsilon$. The first term in Eq. 5 is a confidence-weighted data fidelity loss that aligns $x$ with the diffusion prediction $\tilde{x}$, with deviations weighted according to prediction confidence (i.e., lower weight in regions of high uncertainty). The second term enforces spatial smoothness through total variation regularization, and the third term constrains temporal consistency by penalizing discrepancies between the current roughness or metallic estimate and its backward-warped counterpart.
@@ 3.4 Relighting and Compositing
We perform single-image environment relighting using albedo, depth, normals, roughness and metallic estimated as described above. Shading is computed using Disney’s Principled BRDF [11] using a Cook–Torrance microfacet BRDF with a GGX normal distribution, Smith geometry term, and Schlick Fresnel approximation. Illumination is provided by a high-dynamic-range environment map rendered from the target scene at the approximate actor location. For each environment direction, we evaluate the BRDF contribution modulated by a soft shadow term, which is computed by ray-marching from each surface point along the light direction in view space. At each step, the sample position is projected back into screen space, and the depth buffer is queried to detect occluders. An occlusion factor is accumulated over all steps, then normalized to produce a soft shadow map in the range $[0,1]$.
We composite the gamma-corrected, relit foreground with the target background using the alpha channel from the GOF render.
@@ 4 Experiments
We analyze the performance of diffusion-based methods for relighting across multiple scenes and lighting conditions, and demonstrate that our approach offers advantages in terms of robustness and visual reliability. Our evaluation comprises qualitative results on both real volumetric recordings and synthetic data (Sec. 4.1), complemented by a quantitative analysis on the synthetic dataset (Sec. 4.2). To further demonstrate the effectiveness of individual components of our method, we include an ablation study (Sec. 4.3), and discuss limitations in Sec. 4.4.
For comparison, we benchmark our approach against several recent state-of-the-art methods in neural relighting and rendering, including NVIDIA Diffusion Renderer [43], Neural Gaffer [34], IC-Light [81], Gaussian Opacity Fields (GOF) [77], and Relightable 3D Gaussian (R3DG) [28].
Datasets
Our real-world evaluation is based on an extended set of various scenes from the RIFTCast dataset [85]. Each scene consists of 36 high-resolution color images at a resolution of $24.5$ MP, captured within a calibrated light stage with cameras arranged in a cylindrical rig of 5.25 meters in diameter. We constructed $12$ diverse target environments in Blender to place the captured content.

### v2

@@ Abstract
Figure 1: We integrate diffusion-based decomposition priors with variational methods and image-based lighting to achieve physically plausible, temporally stable relighting of captured volumetric content and seamless compositing into virtual environments.
@@ 1 Introduction
The first challenge, novel view synthesis (NVS), has seen rapid progress in recent years, with methods such as Neural Radiance Fields (NeRF) [39] and 3D Gaussian Splatting [30], along with extensive follow-up work [14], enabling photorealistic rendering from sparse multi-view input and establishing NVS as a reliable foundation for modern volumetric video pipelines. The second challenge, relighting, has been extensively studied, but remains difficult in capture stages where inputs are sparse, lighting is uniform, and material or illumination cues are limited. While one-light-at-a-time (OLAT) capture can enable high-quality relighting [54, 50], it remains impractical for dynamic content, leaving inverse rendering fundamentally underconstrained. Recent data-driven approaches, including diffusion-based decomposition models, offer a promising alternative by directly predicting material and illumination properties [44, 61, 29, 33], but their applicability in volumetric video pipelines remains unclear.
Figure 2: VRAM usage vs. resolution and sequence length. The diffusion renderer exceeds memory limits at moderate resolutions even on an 80 GB GPU, while our method scales efficiently to higher resolutions and longer sequences.
This motivates our central question: Are diffusion relighting models ready for capture-stage compositing? While their progress is promising, it remains unclear whether existing approaches can meet the temporal stability and scalability demands of real-world volumetric capture pipelines: We find that diffusion-based relighting models, whose predictions are inherently stochastic, produce inconsistent material predictions under repeated inference, leading to visible flickering when applied frame by frame. Video-based approaches mitigate this only at the cost of memory requirements that become prohibitive even at moderate resolutions (cf. Fig. 2), necessitating inference on short, overlapping frame chunks that introduce temporal discontinuities.
In this work, we revisit relighting from the perspective of capture-stage compositing and explicitly reformulate diffusion-based relighting as a temporally constrained inference problem. Rather than introducing a new generative diffusion model, we treat state-of-the-art diffusion-based decomposition as a stochastic prior and combine it with explicit temporal regularization and physically motivated rendering. Concretely, we introduce a flow-guided, confidence-aware temporal regularization scheme, together with a lightweight screen-space relighting module inspired by real-time graphics pipelines. By combining learned priors with physically informed design choices, our approach outperforms state-of-the-art relighting methods in temporal stability and visual fidelity, making it suitable for volumetric capture pipelines. In summary, our contributions are
@@
  • An end-to-end volumetric capture relighting pipeline that integrates state-of-the-art image decomposition priors with physically-motivated rendering, enabling consistent and photorealistic results across dynamic sequences.
  • A confidence-aware temporal inference formulation that treats stochastic diffusion predictions as noisy observations and aggregates them into temporally coherent material estimates, substantially reducing flickering.
@@
  • A comprehensive evaluation of diffusion-based relighting approaches in real-world volumetric capture scenarios, highlighting practical trade-offs in quality, temporal stability, and computation cost.
@@ 2 Related Work
aim to recover surface reflectance and illumination for appearance synthesis under novel lighting. Early work formulated this as an ill-posed inverse rendering problem addressed with strong hand-crafted priors [3, 2], including SVBRDF estimation with carefully designed models [1]. With deep learning, these constraints were relaxed, enabling SVBRDF estimation from single images [15] and multi-view observations [16], as well as end-to-end inverse rendering frameworks that jointly recover geometry, reflectance, and illumination using differentiable rendering across single-, dual-, and multi-view settings [56, 31, 32, 47, 53, 7, 13, 36]. Recent years have seen significant progress in neural methods for relighting, which aim to synthesize novel illuminations for existing scenes or objects. Neural Radiance Fields (NeRF) [39] have been extended to support relighting by explicitly modeling material and illumination properties, via intrinsic decomposition of scene properties [5, 63, 46, 49], or via a learned scene representation using One-Light-at-a-Time (OLAT) recordings [54]. The introduction of 3D Gaussian Splatting (3DGS) [30] has led to diverse relighting strategies, via intrinsic decomposition [25, 19, 35, 48, 34], or relighting using OLAT data [17, 4]. Some neural relighting methods avoid 3D reconstruction altogether, operating entirely in image space, e.g. by modelling view- and light-dependent transport as a learned function [62] or training a CNN [55] to predict relit results from sparse lighting samples.
More recently, diffusion models have been applied to relighting due to their strong generative priors. Early works focus on illumination harmonization or background adaptation [44, 61], while others explore relighting via text prompts or explicit light control [29, 60, 59]. Neural Gaffer [29] supports both text- and environment-map–based relighting. While these approaches produce convincing results for single images, Diffusion Renderer [33] extends it to video by unifying inverse and forward rendering, reducing temporal inconsistency at the cost of substantial memory requirements.
@@
aims to generate images of a scene from viewpoints not present in the input observations. Early approaches relied on image-based rendering and depth-guided warping, later extended by learning-based approaches using multiplane images or voxel embeddings [20, 27, 10, 9, 22, 28, 38]. More recently, neural radiance fields became the prevalent approach for novel view synthesis, with subsequent work improving efficiency through alternative scene representations such as low-rank grids and sparse voxels [39, 11, 24]. NeRFs have been extended to dynamic scenes and relightable representations [43, 41, 42, 23, 5, 63, 6, 49], with further work addressing view-dependent appearance, faces, and scene editing [51, 46, 58]. Real-time rendering has likewise become a key focus [40], and recent work has moved toward alternative representations such as 3D Gaussian Splatting [30], which replaces implicit fields with explicit point-based primitives for real-time rendering. Subsequent work has further improved rendering speed and quality [21, 26, 52, 37] and extended Gaussian-based methods beyond basic rendering, including subsurface scattering and real-time relighting [18, 45], as well as surface reconstruction for geometry recovery [57, 12].
@@ 3 Hybrid Relighting Framework
Our method combines three key components as shown in Fig.˜3: volumetric scene representations for novel view synthesis and proxy geometry (Sec.˜3.1), estimation and regularization of intrinsic material properties (Secs.˜3.2 and 3.3) and a physically based relighting stage that produces view-consistent shading under novel illumination (Sec.˜3.4). The final compositing step accounts for indirect interactions such as cast shadows and reflections, ensuring that the subject integrates plausibly with its surroundings. The overall pipeline is designed for offline relighting and compositing and aligns with standard volumetric capture workflows. In the following, we describe each of these stages in detail.
@@ 3.1 Novel View Synthesis & Indirect Interactions
Given a multi-camera capture of a performance, our goal is to render the subject from novel viewpoints and to integrate it into a new environment with plausible indirect light interactions, specifically cast shadows on nearby receivers and reflection contributions, while supporting relighting. To achieve this, we ideally need a NVS representation that provides (i) RGB appearance, (ii) depth maps, (iii) surface normals, and (iv) a proxy geometry suitable for integration into conventional rendering engines. Gaussian Opacity Fields (GOF) [57] are a natural choice, as they jointly optimize for appearance and normals, produce accurate depth estimates, and allow mesh extraction for downstream use.
However, we observe that GOF reconstructions often suffer from artifacts: curved, natural surfaces are approximated as piecewise planar surfaces, and spurious holes may appear. To mitigate these issues, we introduce two additional regularization terms during optimization: a Huber-regularized total variation loss on the rendered depth map, and an edge-aware Laplacian smoothness prior on the Gaussian normal map. For the rendered depth map $z(x,y)$, enforce spatial smoothness via
$\mathcal{L}_{\text{depth}}=\sum_{x,y}H_{\delta}(\tfrac{\partial z}{\partial x})+H_{\delta}(\tfrac{\partial z}{\partial y}),$ (1)
where $H_{\delta}$ denotes Huber loss. To prevent blocky surface approximations while respecting appearance discontinuities, we penalize the Laplacian of the estimated normal field $n(x,y)$ in an edge-aware fashion with respect to the RGB image $I(x,y)$ as
@@
The extracted GOF mesh serves as a proxy mesh that enables indirect light transport effects in conventional rendering engines such as Blender. The proxy mesh is not rendered directly; instead, it is included in the scene only to enable indirect interactions such as shadow casting and reflection contributions, without requiring perfect geometry or material accuracy. To this end, intersections with foreground camera rays are disabled, ensuring that the mesh itself remains invisible in the final image. Since indirect effects are typically low-frequency, coarse geometry suffices for the majority of capture scenarios to render plausible background scenes.
3.2 Stochastic Intrinsic Decomposition
We estimate the unknown G-buffer channels (i.e. roughness and metallic) using learned predictions for view- and light-independent material properties. We consider the problem of predicting roughness $r$ and metallic $\mu$ maps, from a single color image $I\in\mathbb{R}^{H\times W\times 3}$ by employing the intrinsic decomposition model $\mathcal{D}_{\theta}$ from [33] that provides the mapping
$\mathcal{D}_{\theta}:I\rightarrow\{r,\mu\}.$ö (3)
@@ 3.2 Stochastic Intrinsic Decomposition
where $\xi_{i}$ denotes the diffusion noise. These patchwise predictions are then blended into full-resolution maps $\tilde{x}^{(i)}$, yielding an ensemble $\{\tilde{x}^{(1)},\dots,\tilde{x}^{(K)}\}$ from which we compute per-pixel means and variances, thereby suppressing sampling variability and exposing an uncertainty signal useful for downstream processing. We treat the captured RGB as a base color approximating diffuse reflectance under near-uniform capture illumination. While not fully intrinsic albedo, it is stable for relighting and compositing and avoids intrinsic decomposition failures.
@@ 3.3 Temporal Consistency via Optical Flow-Guided Regularization
Next, we address temporally consistent material estimation from stochastic diffusion predictions, treating per-frame diffusion outputs as noisy observations of an underlying time-consistent material field. Let $\tilde{x}\in\{\tilde{r},\tilde{\mu}\}$ denote the per-pixel mean prediction and $\sigma_{x}$ the corresponding normalized standard deviation obtained from multiple diffusion samples. We further define a binary foreground mask $m_{\alpha}=\mathbbm{1}_{\alpha>0.5}$, backward warping operator $\mathcal{W}$ from frame $t+1$ to $t$ based on optical flow, and a validity mask $\mathcal{V}$ that indicates flow vectors which point to valid foreground and are forward/backward consistent. Under this formulation, we estimate temporally consistent material maps by solving the following optimization problem over all frames:
$\underset{x}{\arg\min}\ \sum_{t}\langle H_{\delta}(\tilde{x}_{t}-x_{t}),1-\sigma_{x,t}\rangle+\lambda_{1}TV_{\varepsilon}(x_{t})+\lambda_{2}\left\|\left(\mathcal{W}(x_{t+1})-x_{t}\right)\odot\mathcal{V}\right\|_{2}^{2}$ (5)
where $\odot$ denotes element-wise multiplication, $\langle\cdot,\cdot\rangle$ is the element-wise inner product, $H_{\delta}$ is the (element-wise) Huber loss, and $TV_{\varepsilon}$ the smoothed isotropic total variation for small $\varepsilon$. The objective consists of three terms that jointly enforce data fidelity, spatial smoothness, and temporal consistency, where the first term is a confidence-weighted data fidelity loss that aligns $x$ with the diffusion prediction $\tilde{x}$, with deviations weighted according to prediction confidence (i.e., lower weight in regions of high uncertainty). The second term enforces spatial smoothness through total variation regularization, and the third term constrains temporal consistency by penalizing discrepancies between the current roughness or metallic estimate and its backward-warped counterpart.
@@ 3.4 Relighting and Compositing
We perform single-image environment relighting using base color, depth, normals, roughness and metallic estimated as described above. Shading is computed using Disney’s Principled BRDF [8] using a Cook–Torrance microfacet BRDF with a GGX normal distribution, Smith geometry term, and Schlick Fresnel approximation. Illumination is provided by a high-dynamic-range environment map rendered at the approximate actor location in the target scene. Relighting is performed in screen space: for each pixel, the BRDF is integrated over the full environment map and the resulting illumination is modulated with a soft shadow term computed via view-space ray marching against the depth buffer. The accumulated occlusion yields a shadow factor in $[0,1]$. The gamma-corrected, relit foreground is finally composited over the target background using the alpha channel from the GOF render.
@@ 4 Experiments
We benchmark our approach against recent state-of-the-art methods in neural relighting and rendering, including NVIDIA Diffusion Renderer [33], Neural Gaffer [29], IC-Light [61], and Relightable 3D Gaussian (R3DG) [25]. As reference, we additionally relight our reconstructed proxy mesh using Blender.
Datasets and Implementation Details
Our real-world evaluation is based on various scenes from the RIFTCast dataset [64]. We constructed $12$ diverse target environments in Blender to place the captured content. Furthermore, we created a synthetic dataset of 10 rendered objects in a virtual light stage imitating the above setup. The objects cover diverse geometry and material properties including diffuse and specular reflectance, and are evaluated under 11 HDR environment maps. All experiments are implemented in PyTorch. We use the Adam optimizer with its default parameters.
4.1 Synthetic Analysis of Relighting Accuracy
Table 1: Quantitative comparison of relighting accuracy on synthetic data.
Method Conditioning PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Reference HDR Env. Map $29.42\pm 5.79$ $946\pm 0.040 0.0$$062\pm 0.055$
IC-Light [61] LDR Background $19.41\pm 4.58$ $908\pm 0.081 0.0$$087\pm 0.084$
Neural Gaffer [29] HDR+LDR Env. Map $24.43\pm 4.50$ $917\pm 0.084 0.0$$084\pm 0.092$
R3DG [25] HDR Env. Map $26.15\pm 5.92$ $926\pm 0.067 0.0$$078\pm 0.071$
Diffusion Renderer [33] LDR+Log Env. Map $27.49\pm 6.19$ $937\pm 0.050 0.0$$082\pm 0.071$
Yesnt (Ours) HDR Env. Map $30.19\pm 5.48$ $958\pm 0.028 0.0$$052\pm 0.046$
We quantitatively evaluate relighting accuracy on the synthetic dataset by comparing relit outputs against ground-truth renderings under identical environment maps. Tab.˜1 summarizes the results across standard image-based metrics. Our method consistently achieves the best scores across all three metrics. In particular, it improves PSNR by a clear margin over all competing diffusion-based approaches, while also yielding the highest SSIM and lowest LPIPS. Fig.˜5 further provides qualitative comparisons on the synthetic dataset, showing our relit results alongside ground-truth renderings.
Memory Requirements